OpenStudy (anonymous):

HELP? order from least to greatest. 0.6, -2/3, -1/9, -0.45, -0.4

OpenStudy (lgbasallote):

-2/3, -0.45, -0.4, -1/9, 0.6

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

First off all, using a calculator, find the exact values of the fractions. -2/3 = -0.6666667 -1-9 = -0.1111111 Now, the number furthest from 1 will appear as the first number in the sequence. This is the answer. -2/3, -0.45, -0.4, -1/9, 0.6.

OpenStudy (anonymous):

thanks :D